# Dust in the wind of outbursting young stars

### v1

@@ Abstract
Context. Young Stellar Objects (YSOs) are observed to undergo powerful accretion events known as FU Orionis outbursts (FUors). Such events of episodic accretion are now considered to be common during low mass star formation, wherein the accretion onto the protostar occurs through a surrounding centrifugal disk. Increasing evidence suggests that the magnetic disk winds are crucial for driving disk accretion, as they carry both mass and momentum away from the disk.
@@
Methods. We conduct magnetohydrodynamic (MHD) simulations of formation and evolution of protoplanetary disk (PPD) in the thin-disk limit. We include evolution of dust with two populations and a realistic prescription for viscosity during outbursts, which depends on the local thermal ionization fraction. The disk evolves with the concurrent action of viscosity, self-gravity and magnetic disk winds.
Results. The simulated disk displays outbursting behavior in the early stages, with the duration and frequency of the bursts, their rise times, and brightness amplitudes resembling the observations of FUors. We find that during the outbursts, the winds are over an order of magnitude more dusty, as compared to in quiescence. However, despite this increased dust content, the winds are still dust-depleted as the dust-to-gas ratio is about an order of magnitude lower than the canonical interstellar value of 0.01. The results of our numerical experiments are in general agreement with the available observational findings and they shed a light on the mechanism behind production of dusty winds during outbursting events in YSOs.
@@
Episodic accretion is considered to be an integral part of low-mass star formation (Audard et al., 2014; Fischer et al., 2023). As a young star accretes through the surrounding disk, the system suddenly gets brighter by several magnitudes, and remains in this luminous state for decades or even centuries. These eruptions are named after the prototype FU Orionis system (Herbig, 1977; Hartmann & Kenyon, 1996). These are essentially enhanced accretion events, wherein the disk accretion rate rapidly increases from typically to . Such large and powerful outbursts have significant impact on the mass budget, disk chemistry, ice lines, dust evolution, mineralogy, and dynamics, especially in the innermost few au of the disk. Although several phenomena can possibly explain such outbursts, the mechanism behind episodic accretion remains poorly constrained (Bonnell & Bastien, 1992; Bell & Lin, 1994; Armitage et al., 2001; Vorobyov & Basu, 2005; Bae et al., 2013).
@@
The disk that surrounds a protostar is traditionally considered to be evolving viscously (Shakura & Sunyaev, 1973), with magnetorotational instability (MRI) providing the necessary turbulent viscosity (Balbus & Hawley, 1991; Turner et al., 2014). However, several lines of reasoning indicate that viscosity alone is insufficient to drive the evolution of observed disks (Pinte et al., 2016; Dullemond et al., 2018; Franceschi et al., 2023). Magnetic disk winds have been suggested to play a significant role in the angular momentum transport (Bai, 2016; Pudritz & Ray, 2019; Lesur et al., 2023; Pascucci et al., 2023). The disk winds are distinct from the highly collimated jets often associated with star formation. The winds are essentially magnetocentrifugal in nature and are launched from a large extent of the disk (Blandford & Payne, 1982). These winds are observed as rotating, conical molecular outflows in resolved sources (Güdel et al., 2018; Lee et al., 2021; Launhardt et al., 2023) or low velocity component (LVC) in high spectral resolution observations of forbidden gas lines (Natta et al., 2014; McGinnis et al., 2018), although observationally, they are difficult to distinguish from photoevaporative winds (Alexander et al., 2014; Rab et al., 2023). However, unlike photoevaporative winds, the magnetic winds carry not only mass but also angular momentum from the disk. The effects of magnetic winds on the disk evolution are significant, as the outflow rate is of the order of accretion rate onto the central star (Watson et al., 2016; Tabone et al., 2020).
@@
When it comes to the nature of disk winds during episodic accretion, our understanding is incomplete due to both observational limitations and the complexity of physics involved. Some protostellar jets are observed to show series of shock fronts or knotty structures, which are thought to be causally related to the increased accretion during episodic accretion (Arce et al., 2007; Vorobyov et al., 2018). Although the data is sparse, the evidence suggests that the eruptions are also accompanied by enhanced extended ejections, wherein the mass loss rate of the outflow increases up to an order of magnitude (Calvet et al., 1993; Audard et al., 2014; Milliner et al., 2019). Shearing box MHD simulations suggest that this may be because of the associated increase in the FUV luminosity of the central object (Bai, 2013; Bai et al., 2016). The ionization layer thickness at the disk surface increases with the incident FUV radiation, resulting in an increase in the wind strength (both in terms of mass and angular momentum loss), due to better coupling with the magnetic field. Although it is hard to differentiate outflows from photoevaporative and magnetic winds, both observationally and theoretically, in reality, both of these may coexist as “magneto-thermal disk wind” (Clarke et al., 2005; Bai et al., 2016).
@@
In terms of the dust content, similar to photoevaporative winds, the magnetic disk winds preferentially remove gas and are thus significantly dust depleted (Suzuki et al., 2010; Gorti et al., 2015). Some of the Near Infrared (NIR) interferometric observations of T Tauri stars can be best explained by models of sub-micron-sized dust well above the disk midplane (Labdon et al., 2023). Preliminary observations of spatially resolved disk winds with HST and JWST also point towards their ability transport small dust over large distances (Duchene et al., 2024). Another indicator of dust content in the vicinity of T Tauri stars comes from X-ray observations, which show two component spectra, wherein the photoelectric absorption of the harder component suggests presence of dust depleted gas (Güdel et al., 2007; Güdel, 2008). Although it is thought that sub-micron sized dust dominates the outflows, the distribution of dust within disk winds is not well characterized. In the case of episodic accretion the situation becomes more complex, e.g., because of short- and long-term time variability, projection effects, or interaction with surroundings of the infalling envelope (Clarke et al., 2005; Kóspál et al., 2008). However, recent observations show evidence of fading in young eruptive stars dominated by an increase in circumstellar extinction, which can be interpreted as presence of dusty winds (Nagy et al., 2023). In addition, deep fadings in the brightness of UX Ori-type stars can be explained by eclipses caused by dusty winds with non-axisymmetric distribution or structural inhomogeneity (Shulman & Grinin, 2022; Grinin et al., 2023).
@@
In this paper, we use a sophisticated and comprehensive global model of protoplanetary disk, which contains all the necessary ingredients to investigate the phenomenon of dust ejection during episodic accretion. In Sect. 2 we describe the key components of this model, which are relevant to understanding the subject at hand. In Sect. 3, we elaborate on our results and our findings are summarized in Sect. 4.
@@ 2 Model description
In this paper, we employ the modelling tool set described in (Kadam et al., 2025) and in the interest of brevity, here we summarize the most salient parts. The Eulerian magnetohydrodynamic (MHD) code that is used for modelling the PPDs is called Formation and Evolution Of a Star And its circumstellar Disc (FEOSAD) and the relevant version of this code is described in detail in Vorobyov et al. (2020b). The following description focuses on dust evolution, outburst physics, and magnetic disk winds, as these components are most relevant for understanding the results (see Kadam et al. (2025) for more details). The simulations are conducted in the thin disc limit, wherein the evolution of both gas and dust is considered. The simulations start with the gravitational collapse phase of a starless cloud core, with the initial surface density and rotational profiles consistent with an axisymmetric core collapse (Basu, 1997). The equations of continuity, momentum conservation, and energy transport are solved for the gas component, while the evolution of the magnetic field is considered in the ideal MHD approximation.
@@
The dust in our model consists of two components; submicron-sized small dust, which is fully coupled with the gas, while the grown dust component drifts with respect to the gas. The grown dust exerts a back reaction onto the gas and evolves within the Epstein regime. The small dust can turn into grown dust and vice versa, depending on the dust growth rate and the dynamically varying maximum size of dust grains (Vorobyov et al., 2022). The conversion rate between the dust components is derived based on the assumption that each of the two dust populations have a power-law function that is continuous with a power law index of , and both the minimum size of dust grains and dust mass stay constant during dust growth. If the maximum size of dust grains is reduced, for example, during an outburst due to increased turbulence and temperature, the same governing equations convert grown dust into small dust. Dust growth is limited by dust radial drift, considered self consistently via the solution of dust dynamics equations for grown dust, and by the fragmentation barrier with a fixed fragmentation velocity of 3.0 m s-1.
@@
The magnetorotational instability (MRI) is considered to be the primary driver of disk viscosity. The numerical viscosity is negligible in comparison to the physical one owing to the use of the third-order accurate piecewise-parabolic interpolation scheme (Colella & Woodward, 1984). The disk is allowed to exhibit layered accretion (Gammie, 1996), as the viscous evolution is considered with an adaptive Shakura & Sunyaev (1973) prescription. The thickness of the MRI-active layer is calculated from the ionization fraction, with the assumption that Ohmic diffusivity dominates the disk midplane. The ionization fraction () is obtained by solving the detailed ionization balance equation, which considers collisional ionization and recombination on dust grains as well as due to radiative processes (Dudorov & Khaibrakhmanov, 2014; Vorobyov et al., 2020b). The disk self-gravity takes into account contribution from both the gas and dust components, and the gravitational torque forms a dominant source of mass and angular momentum transport in the early stages (Vorobyov & Basu, 2009).
@@
The outbursts occurring in our simulations resemble the FU Orionis eruptions and are investigated in detail in a series of studies (Kadam et al., 2020; Vorobyov et al., 2020a; Kadam et al., 2022). These outbursts are termed MRI-bursts, since the underlying mechanism relates to sudden ionization of alkali metals in the inner disk due to viscous heating and successive activation of MRI the magnetically dead zone. The thermal ionization considers the contribution from Potassium only, as it dominates other alkali metals and is sufficient for assuming MRI activity (Dudorov & Khaibrakhmanov, 2014; Vorobyov et al., 2020b). The associated value of -parameter during the bursts is assumed to be 0.1, which is motivated by 3D radiation magnetohydrodynamic simulations for FUor disk (Zhu et al., 2020). The disk is considered outbursting above a characteristic threshold value of thermal ionization fraction, (Kadam et al., 2025). In the non-outbursting state, the disk can have , which is consistent with canonical values of protoplanetary disks. In addition, a magnetically dead zone is modeled as weighted average of surface densities of MRI active and dead layers, so that the dead zone corresponds to a lower effective (Bae et al., 2014; Vorobyov et al., 2020b).
@@
The disk winds are magnetocentrifugal in nature, and our model (Kadam et al., 2025) is based on insights gained from shearing box simulations (Bai, 2013). The winds model the extended, conical (non-collimated), slow (low velocity component, LVC) and rotating molecular outflows observed in YSOs. The winds depend on the local properties of the disk (magnetic and gas surface density), as well as its global properties (distance from the central star, stellar mass and luminosity). The winds remove gas mass and angular momentum from the entire extent of the centrifugal disk. The grown dust typically settles to the disk midplane and the winds originate from the disk surface several scale heights above the midplane. Hence the winds are assumed to entrain and remove only the small dust, in proportion to the local small dust-to-gas ratio. Such entrainment of sub-micron-sized dust particles by disk winds is supported by MHD simulations (Tsukamoto et al., 2021; Bino et al., 2022; Rodenkirch & Dullemond, 2022), as well as observations (Labdon et al., 2023; Duchene et al., 2024). During an outburst, the Far-UV (FUV) luminosity of the system increases and the increase in wind strength due to the associated increase in the FUV penetration depth is taken into account.
@@ 3.1 Long-term episodic accretion
The particular simulation discussed henceforth describes evolution of a PPD around a low mass star and is identical to model WI-3a in Kadam et al. (2025). The mass of the initial cloud core is 0.83 , while maximum corresponding to a fully active disk is . The disk wind parameters considered result in a reasonable fit to the observations in terms of global disk properties such as disk mass and radius. Fig. 1 shows the temporal evolution of the disk in terms of characteristic global quantities over the 1 Myr of evolutionary period, starting from the beginning of the cloud core collapse. The Class 0/I boundary (0.036 Myr) is defined as the time when the mass of the star-disk system exceeds half the mass of the initial cloud core, while the Class I/II boundary (0.187 Myr) is considered to be the instance when the envelope accretion rate falls below . The first panel shows the stellar photospheric luminosity and total (stellar+accretion) luminosity, while the second panel shows accretion rate onto the protostar as well as cloud core infall rate onto the disk. The accretion luminosity is punctuated by quasi-periodic and luminous outbursts of approximately 100 , which correspond to accretion rates of as seen in panel b). The accretion at early times is dominated by the outbursts, which continues into the Class II stage. These correspond to MRI-type outbursts, which are triggered in the inner disk by an instability when viscous heating becomes sufficient to thermally ionize the alkali metals, which increases the degree of ionization and hence , causing amplified heating (Bae et al., 2013; Kadam et al., 2020; Vorobyov et al., 2020b). Such outbursts robustly occur in the early stages of the disk evolution, for example, for a wide range of stellar masses (Elbakyan et al., 2021) and low metallicity environments (Kadam et al., 2021).
@@
Panel c) of Fig. 1 shows the gas and dust mass loss rates of the magnetic wind ( and ), integrated over the extent of the centrifugal disk. During quiescence, is approximately equal to the stellar accretion rate, . This is partly achieved in our disk wind model via a multiplicative factor in the mass loss rate (parameter ), which is not well constrained in shearing box simulations (see Kadam et al., 2025, for details). The atomic and molecular lines that trace slower and wider disk winds in the LVC component indicate a similar ratio of on the order of unity (Pascucci et al., 2023). During an outburst, is about one order of magnitude less than . This occurs in the disk wind model through the dependence of the disk wind on the total luminosity of the central object, the latter being a proxy for the FUV irradiation of the disk. Panel d) shows the mass of the star as well as that of the disk. The small discontinuities in both the masses correspond to the episodic accretion, during which a significant amount of mass (approximately a percent of the central star) is accreted onto the star. At 1 Myr, the disk mass is 0.05 , which is almost 10% that of the central star. Thus, at 1 Myr, the system displays properties of a typical Class II TTauri star. We emphasize that the number and magnitude of the outbursts is congruent with observations of outbursting YSOs (e.g., Audard et al., 2014), although the burst statistics are yet to be compared in detail.
@@
The last panel e) of Fig. 1 shows the grown and small dust mass contained within the entire disk. The most salient feature of this plot is that coincident with the outbursts, the grown dust content of the disk suddenly decreases, while the small dust content peaks. Note that all peaks are not captured in the output, as the dust mass content of the disk is derived from 2D output files with a lower temporal resolution. The grown dust is prone to radial drift and thus, it accumulates in the pressure maxima which are formed in the gas rings near the inner edge of the dead zone at a few au from the star (see Kadam et al., 2019, 2022, for details on formation of rings). During an MRI-type burst, the large grains are destroyed due to both increase in temperature and local , since the fragmentation size in inversely proportional to these quantities (Birnstiel et al., 2012). Thus, large dust is converted into small dust, resulting in sudden increase in the small dust content. As the small dust is well-coupled with the gas, it is accreted onto the star as the outburst proceeds, bringing the small dust content near its original baseline.

### v2

@@ Abstract
Context. Young stellar objects (YSOs) have been observed to undergo powerful accretion events known as FU Orionis outbursts (FUors). These types of episodic accretion events are now considered to be commonplace during low-mass star formation, wherein accretion onto the protostar occurs through a surrounding centrifugal disk. Increasing evidence suggests that the magnetic disk winds are crucial for driving disk accretion, as they carry both mass and momentum away from the disk.
@@
Methods. We conducted magnetohydrodynamic (MHD) simulations of the formation and evolution of a protoplanetary disk (PPD) in the thin-disk limit. We included the evolution of dust with two populations and a realistic prescription for viscosity during outbursts, which depends on the local thermal ionization fraction. The disk evolves with the concurrent action of viscosity, self-gravity, and magnetic disk winds.
Results. The simulated disk exhibits outbursting behavior in the early stages, with the duration and frequency of the bursts, their rise times, and brightness amplitudes resembling the properties observed for FUors. We find that during the outbursts, the winds are over an order of magnitude more dusty, as compared to in quiescence. However, despite this increased dust content, the winds are still dust-depleted as the dust-to-gas ratio is about an order of magnitude lower than the canonical interstellar value of 0.01. The results of our numerical experiments are in general agreement with the available observational findings, shedding light on the mechanism behind the production of dusty winds during outbursting events in YSOs.
@@
Episodic accretion is considered to be an integral part of low-mass star formation (Audard et al., 2014; Fischer et al., 2023). As a young star accretes matter via the surrounding disk, the system suddenly gets brighter by several magnitudes and remains in this luminous state for decades or even centuries. These eruptions are named after the prototype FU Orionis system (Herbig, 1977; Hartmann & Kenyon, 1996). These are essentially enhanced accretion events, wherein the disk accretion rate rapidly increases, typically from to . Such large and powerful outbursts have significant impact on the mass budget, disk chemistry, ice lines, dust evolution, mineralogy, and dynamics, especially in the innermost few astronomical units (au) of the disk. Although a number of phenomena could feasibly explain such outbursts, the mechanism behind episodic accretion remains poorly constrained (Bonnell & Bastien, 1992; Bell & Lin, 1994; Armitage et al., 2001; Vorobyov & Basu, 2005; Bae et al., 2013).
@@
The disk that surrounds a protostar is traditionally considered to be evolving viscously (Shakura & Sunyaev, 1973), with magnetorotational instability (MRI) providing the necessary turbulent viscosity (Balbus & Hawley, 1991; Turner et al., 2014). However, several lines of reasoning indicate that viscosity alone is insufficient to drive the evolution of observed disks (Pinte et al., 2016; Dullemond et al., 2018; Franceschi et al., 2023). It has been suggested that magnetic disk winds play a significant role in the angular momentum transport (Bai, 2016; Pudritz & Ray, 2019; Lesur et al., 2023; Pascucci et al., 2023). The disk winds are distinct from the highly collimated jets that are often associated with star formation. The winds are essentially magnetocentrifugal in nature and are launched from a large extent of the disk (Blandford & Payne, 1982). These winds are observed as rotating, conical molecular outflows in resolved sources (Güdel et al., 2018; Lee et al., 2021; Launhardt et al., 2023) or a low-velocity component (LVC) in high-spectral-resolution observations of forbidden gas lines (Natta et al., 2014; McGinnis et al., 2018); however, it is, in fact, observationally challenging to distinguish them from photoevaporative winds (Alexander et al., 2014; Rab et al., 2023). However, unlike photoevaporative winds, the magnetic winds carry not only mass but also angular momentum from the disk. The effects of magnetic winds on the disk evolution are significant, as the outflow rate is of the order of accretion rate onto the central star (Watson et al., 2016; Tabone et al., 2020).
@@
When it comes to the nature of disk winds during episodic accretion, our understanding is incomplete due to both observational limitations and the complexity of physics involved. Some protostellar jets are observed to show series of shock fronts or knotty structures, which are thought to be causally related to the increased accretion during episodic accretion (Arce et al., 2007; Vorobyov et al., 2018). Although the data are sparse, the evidence suggests that the eruptions are also accompanied by enhanced extended ejections, wherein the mass-loss rate of the outflow increases up to an order of magnitude (Calvet et al., 1993; Audard et al., 2014; Milliner et al., 2019). Shearing box MHD simulations suggest that this may be because of the associated increase in the FUV luminosity of the central object (Bai, 2013; Bai et al., 2016). The ionization layer thickness at the disk surface increases with the incident FUV radiation, resulting in an increase in the wind strength (both in terms of mass and angular momentum loss), due to better coupling with the magnetic field. Although it is hard to differentiate outflows from photoevaporative and magnetic winds, both observationally and theoretically, in reality, both of these may coexist as “magneto-thermal disk wind” (Clarke et al., 2005; Bai et al., 2016).
@@
In terms of the dust content, similar to photoevaporative winds, the magnetic disk winds preferentially remove gas and are thus significantly dust depleted (Suzuki et al., 2010; Gorti et al., 2015). Some of the near-infrared (NIR) interferometric observations of T Tauri stars can be best explained by models of sub-micron-sized dust well above the disk midplane (Labdon et al., 2023). Preliminary observations of spatially resolved disk winds with HST and JWST also point towards their ability transport small dust over large distances (Duchene et al., 2024). Another indicator of dust content in the vicinity of T Tauri stars comes from X-ray observations, which show two component spectra, wherein the photoelectric absorption of the harder component suggests presence of dust depleted gas (Güdel et al., 2007; Güdel, 2008). Although it is thought that sub-micron sized dust dominates the outflows, the distribution of dust within disk winds is not well characterized. In the case of episodic accretion, the situation becomes more complex, for instance, because of short- and long-term time variability, projection effects, or interaction with surroundings of the infalling envelope (Clarke et al., 2005; Kóspál et al., 2008). However, recent observations show evidence of fading in young eruptive stars dominated by an increase in circumstellar extinction, which can be interpreted as presence of dusty winds (Nagy et al., 2023). In addition, deep fadings in the brightness of UX Ori-type stars can be explained by eclipses caused by dusty winds with non-axisymmetric distribution or structural inhomogeneity (Shulman & Grinin, 2022; Grinin et al., 2023).
@@
In this paper, we use a sophisticated and comprehensive global model of protoplanetary disk, which contains all the necessary ingredients to investigate the phenomenon of dust ejection during episodic accretion. In Sect. 2, we describe the key components of this model, which are relevant to understanding the subject at hand. In Sect. 3, we elaborate on our results and our findings are summarized in Sect. 4.
@@ 2 Model description
In this paper, we employ the modeling tool set described in (Kadam et al., 2025) and in the interest of brevity, we summarize the most salient parts here. The Eulerian magnetohydrodynamic (MHD) code that is used for modeling a protoplanetary disk (PPD) is called Formation and Evolution Of a Star And its circumstellar Disc (FEOSAD) and the relevant version of this code is described in detail in Vorobyov et al. (2020b). The following description focuses on dust evolution, outburst physics, and magnetic disk winds, as these components are most relevant for understanding the results (see Kadam et al. (2025) for more details). The simulations are conducted in the thin disc limit, wherein the evolution of both gas and dust is considered. The simulations start with the gravitational collapse phase of a starless cloud core, with the initial surface density and rotational profiles consistent with an axisymmetric core collapse (Basu, 1997). The equations of continuity, momentum conservation, and energy transport are solved for the gas component, while the evolution of the magnetic field is considered in the ideal MHD approximation.
@@
The dust in our model consists of two components; submicron-sized small dust, which is fully coupled with the gas, while the grown dust component drifts with respect to the gas. The grown dust exerts a back reaction onto the gas and evolves within the Epstein regime. The small dust can turn into grown dust and vice versa, depending on the dust growth rate and the dynamically varying maximum size of dust grains (Vorobyov et al., 2022). The conversion rate between the dust components is derived based on the assumption that each of the two dust populations have a power-law function that is continuous with a power law index of , and both the minimum size of dust grains and dust mass stay constant during dust growth. If the maximum size of dust grains is reduced, for example, during an outburst due to increased turbulence and temperature, the same governing equations convert grown dust into small dust. Dust growth is limited by dust radial drift, considered self consistently via the solution of dust dynamics equations for grown dust, and by the fragmentation barrier, with a fixed fragmentation velocity of 3.0 m s-1.
@@
The magnetorotational instability (MRI) is considered to be the primary driver of disk viscosity. The numerical viscosity is negligible in comparison to the physical one owing to the use of the third-order accurate piecewise-parabolic interpolation scheme (Colella & Woodward, 1984). The disk is allowed to exhibit layered accretion (Gammie, 1996), as the viscous evolution is considered with an adaptive Shakura & Sunyaev (1973) prescription. The thickness of the MRI-active layer is calculated from the ionization fraction, with the assumption that Ohmic diffusivity dominates the disk midplane. The ionization fraction () is obtained by solving the detailed ionization balance equation, which considers collisional ionization and recombination on dust grains as well as due to radiative processes (Dudorov & Khaibrakhmanov, 2014; Vorobyov et al., 2020b). The disk self-gravity takes into account contribution from both the gas and dust components and the gravitational torque forms a dominant source of mass and angular momentum transport in the early stages (Vorobyov & Basu, 2009).
@@
The outbursts occurring in our simulations resemble the FU Orionis eruptions and are investigated in detail in a series of studies (Kadam et al., 2020; Vorobyov et al., 2020a; Kadam et al., 2022). These outbursts are termed MRI-bursts, since the underlying mechanism relates to sudden ionization of alkali metals in the inner disk due to viscous heating and successive activation of the MRI in a magnetically dead zone. The thermal ionization considers the contribution from Potassium only, as it dominates other alkali metals and is sufficient for assuming MRI activity (Dudorov & Khaibrakhmanov, 2014; Vorobyov et al., 2020b). The associated value of the -parameter during the bursts is assumed to be 0.1, which is motivated by 3D radiation magnetohydrodynamic simulations for FUor disk (Zhu et al., 2020). The disk is considered outbursting above a characteristic threshold value of thermal ionization fraction, (Kadam et al., 2025). In the non-outbursting state, the disk can have a value of , which is consistent with canonical values of protoplanetary disks. In addition, a magnetically dead zone is modeled as weighted average of surface densities of MRI active and dead layers, so that the dead zone corresponds to a lower effective (Bae et al., 2014; Vorobyov et al., 2020b).
@@
The disk winds are magnetocentrifugal in nature and our model (Kadam et al., 2025) is based on insights gained from shearing box simulations (Bai, 2013). The modeled winds exhibit the extended, conical (non-collimated), slow (low velocity component, LVC), and rotating molecular outflows observed in YSOs. The winds depend on the local properties of the disk (magnetic and gas surface density), as well as its global properties (distance from the central star, stellar mass and luminosity). The winds remove the gas mass and angular momentum from the entire extent of the centrifugal disk. The grown dust typically settles to the disk midplane and the winds originate from the disk surface several scale heights above the midplane. Hence, the winds are assumed to entrain and remove only the small dust, in proportion to the low local dust-to-gas ratio. Such an entrainment of sub-micron-sized dust particles by disk winds is supported by MHD simulations (Tsukamoto et al., 2021; Bino et al., 2022; Rodenkirch & Dullemond, 2022), as well as observations (Labdon et al., 2023; Duchene et al., 2024). During an outburst, the Far-UV (FUV) luminosity of the system increases and the increase in wind strength due to the associated increase in the FUV penetration depth is taken into account.
@@ 3.1 Long-term episodic accretion
The particular simulation discussed in this section describes evolution of a PPD around a low-mass star and is identical to model WI-3a in Kadam et al. (2025). The mass of the initial cloud core is 0.83 , while the maximum corresponding to a fully active disk is . The disk wind parameters considered offer a reasonable fit to the observations in terms of global disk properties such as the disk mass and radius. Figure 1 shows the temporal evolution of the disk in terms of the characteristic global quantities over the 1 Myr of evolutionary period, starting from the beginning of the cloud core collapse. The Class 0/I boundary (0.036 Myr) is defined as the time when the mass of the star-disk system exceeds half of the mass of the initial cloud core, while the Class I/II boundary (0.187 Myr) is considered to be the point at which the envelope accretion rate falls below . The first panel shows the stellar photospheric luminosity and total (stellar+accretion) luminosity, while the second panel shows accretion rate onto the protostar as well as the cloud core infall rate onto the disk. The accretion luminosity is punctuated by quasi-periodic and luminous outbursts of approximately 100 , which correspond to accretion rates of as seen in panel b). The accretion at early times is dominated by the outbursts, which continues into the Class II stage. These correspond to MRI-type outbursts, which are triggered in the inner disk by an instability when viscous heating becomes sufficient to thermally ionize the alkali metals, which increases the degree of ionization – and, hence, increasing and causing amplified heating (Bae et al., 2013; Kadam et al., 2020; Vorobyov et al., 2020b). Such outbursts robustly occur in the early stages of the disk evolution; for example, for a wide range of stellar masses (Elbakyan et al., 2021) and low metallicity environments (Kadam et al., 2021).
@@
Panel c in Fig. 1 shows the gas and dust mass-loss rates of the magnetic wind ( and ), integrated over the extent of the centrifugal disk. During quiescence, is approximately equal to the stellar accretion rate, . This is partly achieved in our disk wind model via a multiplicative factor in the mass-loss rate (parameter ), which is not well constrained in shearing box simulations (see Kadam et al., 2025, for details). The atomic and molecular lines that trace slower and wider disk winds in the LVC component indicate a similar ratio of on the order of unity (Pascucci et al., 2023). During an outburst, is about one order of magnitude less than . This occurs in the disk wind model through the dependence of the disk wind on the total luminosity of the central object, the latter being a proxy for the FUV irradiation of the disk. Panel d shows the mass of the star as well as that of the disk. The small discontinuities in both the masses correspond to the episodic accretion, when a significant amount of mass (approximately a percent of the central star) is accreted onto the star. At 1 Myr, the disk mass is 0.05 , which is almost 10% that of the central star. Thus, at 1 Myr, the system displays properties of a typical Class II TTauri star. We emphasize that the number and magnitude of the outbursts is congruent with observations of outbursting YSOs (e.g., Audard et al., 2014), although the burst statistics have yet to be compared in detail.
@@
The last panel e of Fig. 1 shows the grown and small dust mass contained within the entire disk. The most salient feature of this plot is that (coincident with the outbursts) the grown dust content of the disk suddenly decreases, while the small dust content peaks. We note that all peaks are not captured in the output, as the dust mass content of the disk is derived from 2D output files with a lower temporal resolution. The grown dust is prone to radial drift and, thus, it accumulates in the pressure maxima which are formed in the gas rings near the inner edge of the dead zone at a few au from the star (see Kadam et al., 2019, 2022, for details on formation of rings). During an MRI-type burst, the large grains are destroyed due to both increase in temperature and local , since the fragmentation size in inversely proportional to these quantities (Birnstiel et al., 2012). Thus, large dust is converted into small dust, resulting in sudden increase in the small dust content. As the small dust component is well-coupled with the gas, it is accreted onto the star as the outburst proceeds, bringing the small dust content close to its original baseline.